Which of the following statements is true?
A)i^99 B)i^0+i^18=0 C)1/i^3=-i D)(i^3)^5=1

Answer by fractalier(6550)

Choice B)i^0+i^18=0
because i^0 = 1 and i^18 = -1 and 1 - 1 = 0